# Present address: ]Lehrstuhl für Theoretische Chemie, Ruhr-Universität Bochum, 44780 Bochum, Germany

### v3

@@ II Results and Discussion
Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence . The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium . Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Figure S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth. Our findings also have important implications for organic molecules in interstellar space. The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic . It is unclear how those complex organic species are generated. Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular . Earth’s or other giant planets’ interiors also provide a protective environment that filters out harmful radiation and buffers against drastic physico-chemical variations.
@@ III Conclusions
Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. Particularly, different from that the six-membered ring form of ribose is the majority in solutions at ambient conditions, it was found that the formation of the five-membered ring form of ribose is more favored at extreme P-T conditions, which provides the necessary component for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life. The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions. Our comprehensive studies present a novel perspective extending the cradle of life in planetary interiors. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
Methods
@@ III.1 Ab initio molecular dynamics simulations (AIMD)
We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture . We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized . It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe . We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations. The simulation boxes have periodic boundary conditions. We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave kinetic energy cutoff to 85 Ry and the cutoff for stress confinement potential to 70 Ry focher1994structural . We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave kinetic energy cutoff was 65 Ry. During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Figure S2). Our total AIMD simulation time exceeds 2.5 ns.
@@ Acknowledgements
This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310), and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083). R.T. would like to acknowledge support from the Original Exploration Plan-Key project (42150104). Contributions by D.A.S. constitute material based upon work supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Geosciences program under Award Number DE-SC0019830 as well as NSF Petrology and Geochemistry Grant Number 2032039. Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China.

### v4

@@ II Results and Discussion
Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence . The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some forms of plate tectonics may have already been active around 4.0 billion years ago, and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium . Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively (see Fig. S16)holm2009reduction , and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth. Our findings also have important implications for the synthesis of organic molecules in interstellar space. The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic . The formation mechanism of those complex organic species may be intricate. Our study suggests that the extreme P and T inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, could be considered as one important condition to help to synthesize some of those complex organic molecules postberg2018macromolecular . Earth’s or other giant planets’ interiors also provide a protective environment that filters out harmful radiation and buffers against drastic physico-chemical variations.
Although our simulations were conducted under extreme P-T conditions, our study may also have implications for experiments at lower pressures and temperatures. We found that high pressures can facilitate the synthesis and stabilization of large biomolecules, aligning with some experimental studies conducted at lower pressures otake2011stability ; ohara2007pressure ; furukawa2012abiotic ; pedreira2019spontaneous . It’s important to note that some experiments were performed under water-poor conditions otake2011stability ; ohara2007pressure ; furukawa2012abiotic , whereas we discovered that supercritical water plays a crucial role in the formation of biomolecules.
@@ III Conclusions
Using the molecules formed, we have further explored the formation pathways of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. Particularly, different from that the six-membered ring form of ribose is the majority in solutions at ambient conditions, it was found that the formation of the five-membered ring form of ribose is more favored at extreme P-T conditions, which provides the necessary component for the generation of RNA molecules. Moreover, these compounds also have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origin of life. The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions. Our comprehensive studies underscore the profound impact of extreme P-T conditions on the abiotic synthesis of organic matter and provide valuable insights into the mechanisms underlying the formation of crucial biomolecules for building blocks of life, which presents a novel perspective extending the cradle of life in planetary interiors.
Computational Methodology
@@ III.1 Ab initio molecular dynamics simulations (AIMD)
We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture . We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized . It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe . We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations. The simulation boxes have periodic boundary conditions. We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave kinetic energy cutoff to 85 Ry and the cutoff for stress confinement potential to 70 Ry focher1994structural . We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave kinetic energy cutoff was 65 Ry. During NVT simulations, we selected random time points to carry out independent simulations with the cutoff of 85 Ry to monitor pressure (Fig. S2). For each P-T point, the NVT simulation time exceeds 400 ps (Fig. S3). Our total AIMD simulation time exceeds 2.5 ns.
In our AIMD simulations, we calculated the forces on the nuclei using quantum mechanics, while treating nuclei including protons as classical particles. If we consider quantum nuclear effects, we anticipate an increase in the proton transfer rate, resulting in more active reactions involving proton transfer. However, at high temperatures, the significance of quantum effects may diminish compared to ambient conditions ceriotti2016nuclear .
@@
In our simulations, the calculated free energies are converged within 0.2 kcal/mol in the last 10 ps. The test of convergence of free energy calculations is shown in Fig. S17.
Supporting information
Details of the simulations, Mole percents of C-N containing species over time, hydrocarbon products, free energy calculations, possible synthesis of ribose and uracil and their ring structures, compositions of fluids as a function of oxygen fugacity, test of convergence of free energy calculations.
@@ Acknowledgements
This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310). R.T. would like to acknowledge support from the Original Exploration Plan-Key project (42150104). Contributions by D.A.S. constitute material based upon work supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Geosciences program under Award Number DE-SC0019830 as well as NSF Petrology and Geochemistry Grant Number 2032039. Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China.